Please help!!!!!!
1) (3x+14)^2+25=0
2) 25x^2=13
3) 2(x-5)^2=3
4) x^2+4x+4=7
Thanks!!!!!!

Answer by stanbon(75887)

Please help!!!!!!
1) (3x+14)^2+25=0
FYI: a^2+b^2=(a+bi)(a-bi)
Your Problem:
[3x+14+5i][3x+14-5i]
x=(1/3)(-145-5i) or x=(1/3)(-14+5i)
==========================
2) 25x^2=13
x^2=13/25
x=(sqrt13)/5 or x=-(sqrt13)/5
============================
3) 2(x-5)^2=3
(x-5)^2=3/2
(x-5)=sqrt(3/2) or x-5=-(sqrt(3/2))
x=5+(sqrt6)/6 or x=5-[(sqrt6)/6]
=============================
4) x^2+4x+4=7
(x+2)^2=7
(x+2)=sqrt7 or (x+2)=-sqrt7
x=-2+sqrt7 or x=-2-sqrt7
==============
Cheers,
Stan H.